# RLSA-PFL: Robust Lightweight Secure Aggregation with Model Inconsistency Detection in Privacy-Preserving Federated Learning

### v2

@@
Abstract
Federated Learning (FL) allows users to collaboratively train a global machine learning model by sharing local model only, without exposing their private data to a central server. This distributed learning is particularly appealing in scenarios where data privacy is crucial, and it has garnered substantial attention from both industry and academia. However, studies have revealed privacy vulnerabilities in FL, where adversaries can potentially infer sensitive information from the shared model parameters. In this paper, we present an efficient masking-based secure aggregation scheme utilizing lightweight cryptographic primitives to mitigate privacy risks. Our scheme offers several advantages over existing methods. First, it requires only a single setup phase for the entire FL training session, significantly reducing communication overhead. Second, it minimizes user-side overhead by eliminating the need for user-to-user interactions, utilizing an intermediate server layer and a lightweight key negotiation method. Third, the scheme is highly resilient to user dropouts, and the users can join at any FL round. Fourth, it can detect and defend against malicious server activities, including recently discovered model inconsistency attacks. Finally, our scheme ensures security in both semi-honest and malicious settings. We provide security analysis to formally prove the robustness of our approach. Furthermore, we implemented an end-to-end prototype of our scheme [^1] The fully foundational end-to-end source code will be released upon publication.. We conducted comprehensive experiments and comparisons, which show that it outperforms existing solutions in terms of communication and computation overhead, functionality, and security.
Index Terms:
@@ Abstract
Federated Learning, Machine Learning, Privacy, Secure Aggregation, User Dropouts, Model Inconsistency Attack.
1 Introduction
Federated Learning (FL) is a distributed machine learning (ML) approach designed to enhance user privacy during the model training process. Instead of sending user data to a central server, FL only shares local model updates after training happens on the users’ devices. The central server then combines these updates to create a global model without ever accessing the actual data [1]. This distributed system reduces the chance of sensitive information being exposed, improving both privacy and security. FL is already being used in areas like healthcare [2], telecommunications [3], the Internet of Things (IoT) [4], and smart cities [5].
However, FL on its own cannot fully eliminate privacy risks. Zhu et al. [6] showed that attackers, including the central server, could still reconstruct users’ local datasets by exploiting the gradients shared during model updates. Similar risks have been found in other studies [7, 8, 9], where researchers pointed out that sharing gradients with the central server could reveal sensitive user data. As a result, additional security measures are needed to fully protect privacy in FL. One popular solution is combining FL with privacy-preserving techniques like differential privacy (DP) [1]. However, DP comes with a trade-off: the stronger the privacy protection, the more it deteriorates the utility of the global model [10].
Secure aggregation is an alternative promising privacy-preserving technique that has gained significant attention for its advantages over DP, such as maintaining accuracy, simplicity in implementation, and avoiding trade-offs in utility [11]. In this method, users cryptographically hide their local model updates before sharing them with the central server. The server then aggregates these hidden updates without being able to view the individual contributions. This ensures that no single party, including the central server, can reconstruct individual model parameters while still obtaining an accurate aggregated global model. Secure aggregation effectively mitigates privacy risks like gradient leakage, which could otherwise expose sensitive data [1].
In general, secure aggregation relies on cryptographic primitives such as homomorphic encryption (HE), secure multi-party computation (SMC), and masking-based techniques [1]. Among these, masking-based secure aggregation is often considered more practical than the others due to its relatively lower computational and communication costs [11]. Several masking-based secure aggregation protocols have been proposed in recent years, including foundational techniques like the BBGLR protocol by Bonawitz et al. in [12]. BBGLR employs a cryptographic masking technique that allows users to securely mask their local model updates before aggregation, ensuring that the central server cannot reconstruct individual updates while still generating the global model. Building on this work, various protocols (e.g., [13, 14, 15, 16, 17, 18]) have introduced improvements to enhance security and reduce computational or communication overhead, optimizing FL for diverse real-world applications.
However, the BBGLR protocol and its variants have some limitations. One issue is the computational overhead for users due to the user-to-user secret-sharing mechanism. Users need to generate shared keys with their neighbors and perform cryptographic operations like additive secret sharing, which increases computational complexity, especially for devices with limited resources. Another challenge is the communication overhead, as users must exchange secret shares with neighbors in every round of training, requiring a fresh setup each time [19]. This imposes a significant communication burden, particularly in large-scale FL systems. Since secret sharing is repeated in each round, the scalability becomes a concern, as the overhead grows with both the number of participants and the training rounds. This problem is even more pronounced in scenarios with many users and rounds, as noted in [19].
Moreover, a recent study by Pasquini et al. [20] demonstrated that the BBGLR protocol and it’s variants are vulnerable to a new type of attack known as model inconsistency, particularly when the central server is malicious. This issue also affects other FL systems using secure aggregation techniques like SMC and HE. The key idea behind this attack is that the server controls the updates used in secure aggregation. A malicious server can manipulate these updates to make the final results reveal information about one or more targeted users, undermining the protection secure aggregation is supposed to offer when the central server is compromised.
1.1 Our Contribution
In this paper, we propose a lightweight masking-based secure aggregation scheme for FL that can also detect model inconsistency attacks caused by a malicious server. Our scheme is lightweight in the sense that it only requires one setup phase for entire rounds of FL training, unlike the BBGLR protocol, which requires repeated setups. This reduces both the communication and computational load on users. Additionally, our scheme utilizes a lightweight additive secret-sharing mechanism with key negation for masking, which is more computationally efficient than the BBGLR protocol’s use of the comparatively expensive Shamir’s Secret Sharing.
Further, our scheme uses a small number of intermediate servers instead of having users interact with each other. These servers collect masked models from the users and assist the central server, referred to as the Aggregator, in handling user dropouts. Adding this middle layer offers two key benefits. First, it removes the need for user-to-user interaction, which would otherwise impose an unacceptable burden on resource-limited users [17]. Second, intermediate servers are generally more trusted and reliable than other users, as they could be service providers or entities with higher reputations, resources, and trustworthiness. For example, in real-world deployments, companies like Amazon Web Services (AWS), and Microsoft Azure could act as intermediate servers, providing the necessary infrastructure to handle aggregation, ensuring trustworthiness, and maintaining high performance for scalability.
Our scheme also includes an efficient global model consistency detection mechanism, though it comes with some extra communication costs. We use a message authentication code (MAC) to check for any inconsistencies in the global model received by each user. The intermediate servers play a role in receiving the MACed global models. As long as at least one intermediate server and two users are trustworthy, our scheme can successfully detect model inconsistency attacks.
In summary, our scheme offers the following contributions:
  • We propose a novel secure aggregation method using a lightweight additive secret-sharing mechanism with key negation. It is highly resilient to user dropouts without requiring users to interact with each other or go through multiple complicated setup phases.
  • Our scheme introduces an efficient mechanism to detect model inconsistency attacks, with minimal communication costs.
  • We provide formal security proofs demonstrating our scheme’s robustness against both semi-honest and malicious adversaries.
  • We thoroughly analyze and compare our scheme with closely related works, showing superior performance across computation, communication, functionality, and security.
  • We implemented our scheme in an end-to-end manner and provided comprehensive experimental results, demonstrating its practicality and effectiveness.
The organization of this paper is as follows: Section 2 provides a brief overview of related works in privacy-preserving aggregation techniques in FL. The cryptographic primitives used in our scheme are presented in Section 3. Section 4 outlines the system architecture, threat model, assumptions, and goals of our scheme. Section 5 details our proposed scheme. The security analysis and performance evaluation are covered in Sections 6 and Section 7, respectively. Finally, the paper concludes in Section 8.
TABLE I: Comparison of Computation Complexity, Communication Overhead, Security, and Functionality between Our Scheme and Notable Masking-Based Secure Aggregation Schemes.
Schemes Computation Overhead Communication Overhead Threat Model Dropout Communication Round No. of Setup Phase Model Incon. Attack Detection
@@ 1.1 Our Contribution
SecAgg [12] $\mathcal{O}(m^{2}+m\cdot|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$ $\mathcal{O}(m+|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$ Malicious $\frac{m}{3}$ $4$ $|i|$ No
SecAgg+ [13] $\mathcal{O}(|v|\cdot\log m+\log m)$ NA $\mathcal{O}(m\cdot|v|+m\log m)$ $\mathcal{O}(|v|+\log m)$ NA $\mathcal{O}(m\cdot|v|+m\log m)$ Malicious $\sigma\cdot m$ $3$ $|i|$ No
FastSecAgg [21] $\mathcal{O}(|v|\log m)$ NA $\mathcal{O}(|v|\cdot\log m)$ $\mathcal{O}(m+|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$ Semi Honest $\frac{m}{2}-1$ $3$ $|i|$ No
EDRAgg [17] $\mathcal{O}(m^{2}+|v|)$ NA $\mathcal{O}(m+|v|)$ $\mathcal{O}(m+|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$ Malicious $\frac{m}{2}-1$ $3$ $|i|$ No
Flamingo [19] $\mathcal{O}(S+d)$ $\mathcal{O}(d^{2}+\delta\cdot S\cdot m+(1-\delta)m+\epsilon\cdot m^{2})$ $\mathcal{O}(m|v|+m^{2})$ $\mathcal{O}(m+|v|)$ $\mathcal{O}(d+\delta\cdot S\cdot m+(1-\delta)m)$ $\mathcal{O}(m(|v|+d+S))$ Malicious $\frac{\delta_{d}+\eta_{d}}{3}$ $3$ $1$ Yes
Our Scheme $\mathcal{O}(|v|)$ $\mathcal{O}(m+|v|)$ $\mathcal{O}(m+|v|)$ $\mathcal{O}(|v|)$ $\mathcal{O}(m\cdot|v|)$ $\mathcal{O}(m\cdot|v|)$ Malicious $m-2$ $2$ $1$ Yes
@@
The set of decryptors in Flamingo represented as the intermediate servers for comparison; $m$ represents the total number of participating users; $|v|$ represents the size of the input vector; NA: Not applicable; $\sigma$ represents a security parameter and $\sigma\cdot m\leq\frac{m}{3}$; $|i|$ represents the total number of rounds in the training phase; $d$ represents the number of decryptors; $\delta$ represents user dropout rate; $S$ represents the upper bound on the number of neighbors of a user; $\eta_{d}$ represents the number of corrupted decryptors; $\delta_{d}$ represents decryptor dropout rate; $\epsilon$ represents the graph generation parameter in [19].
2 Related Work
@@
Secure aggregation in FL protects individual users’ local model parameters from being disclosed to the central aggregator server. This mechanism is also referred to as privacy-preserving aggregation [1]. Differential privacy (DP), homomorphic encryption (HE), secure multi-party computation (SMC), and masking methods are being used for privacy-preserving secure aggregation in FL [11], [1]. Our proposed scheme falls within the masking-based method. Therefore, we briefly introduce the other secure aggregation methods while providing a detailed discussion of the masking-based approach.
2.1 DP, HE, and SMC Based Schemes

### v3

# Setup Once, Secure Always: A Single-Setup Secure Federated Learning Aggregation Protocol with Forward and Backward Secrecy for Dynamic Users

@@
Abstract.
Federated Learning (FL) enables multiple users to collaboratively train a machine learning model without sharing raw data, making it suitable for privacy-sensitive applications. However, local model or weight updates can still leak sensitive information. Secure aggregation protocols mitigate this risk by ensuring that only the aggregated updates are revealed. Among these, single-setup protocols, where key generation and exchange occur only once, are the most efficient due to reduced communication and computation overhead. However, existing single-setup protocols often lack support for dynamic user participation and do not provide strong privacy guarantees such as forward and backward secrecy.
In this paper, we propose a new secure aggregation protocol that requires only one setup operation for the entire FL training and allows new users to join or leave at any round. It employs lightweight symmetric homomorphic encryption with a key negation technique to efficiently mask updates, without user-to-user communication— unlike the existing protocols. To defend against model inconsistency attacks, we introduce a simple verification mechanism using message authentication codes (MACs). Our protocol is the first to combine forward/backward secrecy, dropout resilience, and model integrity verification in a single-setup design. We provide formal security proofs and implement an end-to-end prototype, which source code has been released [^1] Our source code is available at https://anonymous.4open.science/r/scatesfl-44E1 for performance testing, and at https://anonymous.4open.science/r/scatesfl-815A for real-world accuracy evaluation.. Our experimental results show that our protocol reduces user-side computation by approximately $99\%$ compared to state-of-the-art protocols like e-SeaFL (ACSAC’24), making it highly practical for real-world FL deployments, especially on resource-constrained devices.
@@ Abstract.
Privacy, Secure Aggregation, Federated Learning, Model Inconsistency, Forward Secrecy, Backward Secrecy, Machine Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Computing methodologies Machine learning
1. Introduction
Recent research in Federated Learning (FL) has revealed serious privacy concerns. Studies such as (Zhu et al., 2019; Geiping et al., 2020; Zhou et al., 2023; Jeon et al., 2024) demonstrate that adversaries, including a potentially malicious central server, can reveal private information of local datasets by analyzing the gradients exchanged during model updates. In response to these threats, secure aggregation has gained significant attention (Chen et al., 2024). This cryptographic technique safeguards user privacy by preventing the server from accessing individual model updates, thereby mitigating the risk of gradient-based reconstruction attacks (Yin et al., 2021).
Masking-based secure aggregation technique has recently garnered significant attention from the research community due to its efficiency, scalability, and strong privacy guarantees in federated learning settings (Chen et al., 2024). The core idea behind this technique is to conceal users’ local model updates using cryptographic masks, which are canceled out during the aggregation phase at the server (et al., 2017). This ensures that the server can compute only the aggregated result without ever accessing any individual update. In their seminal work, Bonawitz et al. (et al., 2017) introduced the first practical and scalable masking-based secure aggregation protocol, which tolerates user dropouts and minimizes user-side overhead. Building on this foundation, a range of protocols have been developed to enhance security and functionality while reducing communication and computation overhead (Bell et al., 2020; Eltaras et al., 2023; et al., 2021a; Guo et al., 2021; Liu et al., 2023; Xu et al., 2020).
However, Bonawitz et al. (et al., 2017) protocol requires a setup operation in each training round, during which users establish pairwise cryptographic keys. This setup phase is similarly adopted by many subsequent protocols (Bell et al., 2020; Eltaras et al., 2023; et al., 2021a; Guo et al., 2021; Liu et al., 2023; Xu et al., 2020). While this design enhances resilience to user dropout and preserves privacy, it introduces additional communication and computation overhead. As noted in (Ma et al., 2023), these costs become particularly significant in large-scale FL deployments, rendering such approaches less practical for real-world applications.
To address the limitations of Bonawitz et al. (et al., 2017) and its variants (Bell et al., 2020; Eltaras et al., 2023; et al., 2021a; Guo et al., 2021; Liu et al., 2023; Xu et al., 2020), Ma et al. (Ma et al., 2023) (S&P’23) and Behnia et al. (Behnia et al., 2024) (ACSAC’24) proposed two masking-based secure aggregation protocols, named Flamingo and e-SeaFL, respectively. Both protocols utilize a single setup operation that supports multiple training rounds in FL. This design simplifies the training process by reducing communication and computation costs, thereby accelerating overall convergence. The core idea is to pre-establish pairwise secrets or seed materials between users in Flamingo (Ma et al., 2023) and between users and assistant nodes in e-SeaFL (Behnia et al., 2024) during the initial setup. These long-term seeds or secrets are then used to deterministically derive per-round masking values using lightweight cryptographic primitives such as pseudo-random functions (PRFs), effectively eliminating the need for costly setup phases in each round.
However, several challenges remain with Flamingo and e-SeaFL. Most critically, these protocols lack forward and backward secrecy. Although they eliminate the need for repeated setup in each round by using long-term pairwise secrets, this design introduces a vulnerability: if a long-term secret key is compromised, an adversary could reconstruct all past and future masking values for that user within the same training session. This concern is particularly relevant in FL settings, where training can span days and devices may operate in untrusted or adversarial environments. Consequently, forward and backward secrecy remain a crucial yet unresolved requirement in these single-setup aggregation protocols. We provide a detailed analysis of forward and backward secrecy for e-SeaFL and Flamingo in Appendix A. Second, these protocols do not support dynamic participation; users must be registered before the initial setup phase and cannot join the training session once it has started. This limitation reduces their practicality in real-world FL deployments, where user availability is often unpredictable. Finally, handling user dropout remains inefficient with Flamingo. It can tolerate up to one-third of corrupted users, including both regular and decryptor users.
In this paper, we introduce a lightweight masking protocol for FL based on additive symmetric homomorphic encryption (HE) with key negation. Like Flamingo and e-SeaFL, our protocol requires only a single setup operation for all rest FL training rounds, significantly reducing communication overhead. However, our design offers key improvements in both security and efficiency. Unlike Flamingo and e-SeaFL, which rely on PRF operations using long-term secret seeds to generate masking values for each element of the model vector, our protocol uses simple random secrets directly as masking parameters. This eliminates the need for per-element PRF evaluations, resulting in substantially lower computational cost and making our approach highly efficient and practical— especially for resource-constrained devices. To further enhance scalability and robustness, our protocol introduces a small set of intermediate servers or edge servers (which may be selected users or assistant nodes similar to Flamingo and e-SeaFL, respectively) to assist in partial aggregation. These intermediate servers collect masked model updates from users and support the central server, referred to as the Aggregator, in managing user dropouts. This intermediate layer provides three key advantages:
  • First, it eliminates the need for user-to-user interaction unlike the Bonawitz et al. (et al., 2017) and its variants (Bell et al., 2020; Eltaras et al., 2023; et al., 2021a; Guo et al., 2021; Liu et al., 2023; Xu et al., 2020), which would otherwise impose a significant burden on resource-constrained users (Liu et al., 2023).
  • Second, avoiding direct user interactions reduces the trust assumptions required in existing protocols, such as those proposed by Bonawitz et al. (et al., 2017) and its variants (Bell et al., 2020; Eltaras et al., 2023; et al., 2021a; Guo et al., 2021; Liu et al., 2023; Xu et al., 2020), for maintaining privacy. This design choice also mitigates the risk of user collusion. Our protocol maintains privacy guarantees as long as at least two users behave honestly, a minimal assumption in FL, ensuring that their individual model updates remain protected.
  • Third, intermediate servers in hierarchical FL architectures offer a practical and scalable solution for handling aggregation and privacy-preserving operations. As demonstrated by Ratnayake et al. (Ratnayake et al., 2025), these servers are typically deployed on infrastructure with sufficient computational capacity and reliable connectivity, such as edge or fog nodes, making them well-suited for tasks like model aggregation, homomorphic encryption, and differential privacy noise addition.
Notably, the inclusion of intermediate servers in our protocol aligns with the trust model adopted in several recent state-of-the-art works (Wang et al., 2018; Kairouz et al., 2021; Ma et al., 2023), where helper nodes or designated users assist the aggregator during the aggregation phase. More recently, Wang et al. (Wang et al., 2025) introduced a similar approach involving edge or intermediate servers. Consistent with these models, our protocol remains robust against collusion attacks, provided that at least one intermediate server behaves honestly. Overall, our approach enhances both performance and security while minimizing the trust assumptions placed on users, without increasing the trust requirements for intermediate servers.
Contributions
Our work makes the following key contributions:
  • We propose a new secure aggregation protocol for FL that uses lightweight symmetric homomorphic encryption with a key negation technique. It requires only one setup for all training rounds and avoids costly per-element PRF operations, making it more efficient than existing state-of-the-art protocols like Flamingo and e-SeaFL.
  • Our protocol is the first single-setup secure aggregation protocol to support dynamic user participation, allowing users to join or leave at any round without reinitialization or user-to-user communication— an important feature missing in existing single-setup protocols like Flamingo and e-SeaFL. It is also the first single-setup secure aggregation protocol to achieve both forward and backward secrecy, ensuring that compromise of long-term keys does not reveal past or future model updates.
  • We introduce a lightweight model verification method to detect model inconsistency attacks using Message Authentication Codes (MACs) and hash functions. This approach avoids expensive cryptographic operations and adds an essential layer of protection for secure aggregation protocols in FL (Pasquini et al., 2022).
  • We provide formal security proofs under both semi-honest and malicious adversarial models, demonstrating robustness against collusion and active attacks. We also present a comprehensive comparison with the most efficient state-of-the-art protocol e-SeaFL, showing that our protocol achieves better performance in terms of computation, communication, security, and functionality. We implement a fully end-to-end prototype and release open-source code with extensive experimental results: https://anonymous.4open.science/r/scatesfl-44E1 (performance testing) and https://anonymous.4open.science/r/scatesfl-815A (real-world accuracy evaluation).
Table 1. Comparison of Functionality and Security
Threat Model Forward & Backward Secrecy Communication Round Setup Round(s) User Dropout Model Verification Dynamic User
SecAgg (et al., 2017) Malicious yes 4 $|i|$ $\frac{m}{3}$ no no
SecAgg+ (Bell et al., 2020) Malicious yes 3 $|i|$ $\sigma\cdot m$ no no
FastSecAgg (Kadhe et al., 2020) Semi Honest yes 3 $|i|$ $\frac{m}{2}-1$ no no
EDRAgg (Liu et al., 2023) Malicious yes 3 $|i|$ $\frac{m}{2}-1$ no no
Flamingo (Ma et al., 2023) Malicious no 3 1 $\frac{\delta_{d}+\eta_{d}}{3}$ yes no
e-SeaFL (Behnia et al., 2024) Malicious no 2 1 m2m-2 yes no
Ours Malicious yes 2 1 m2m-2 yes yes
$m$ represents the total number of participating users; $\sigma$ represents a security parameter and $\sigma\cdot m\leq\frac{m}{3}$; $|i|$ represents the total number of rounds in the training phase; $\eta_{d}$ represents the number of corrupted decryptors; $\delta_{d}$ represents decryptor dropout rate.
Table 2. Asymptotic Computation and Communication Overhead Comparison
Protocols Computation Overhead Communication Overhead
@@ 1. Introduction
SecAgg (et al., 2017) $\mathcal{O}(m^{2}+m\cdot|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$ $\mathcal{O}(m+|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$
SecAgg+ (Bell et al., 2020) $\mathcal{O}(|v|\cdot\log m+\log m)$ NA $\mathcal{O}(m\cdot|v|+m\log m)$ $\mathcal{O}(|v|+\log m)$ NA $\mathcal{O}(m\cdot|v|+m\log m)$
FastSecAgg (Kadhe et al., 2020) $\mathcal{O}(|v|\log m)$ NA $\mathcal{O}(|v|\cdot\log m)$ $\mathcal{O}(m+|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$
EDRAgg (Liu et al., 2023) $\mathcal{O}(m^{2}+|v|)$ NA $\mathcal{O}(m+|v|)$ $\mathcal{O}(m+|v|)$ NA $\mathcal{O}(m^{2}+m\cdot|v|)$
Flamingo (Ma et al., 2023) $\mathcal{O}(S\cdot|v|+d\cdot|v|)$ $\mathcal{O}(d^{2}+\delta\cdot S\cdot m+(1-\delta)m+\epsilon\cdot m^{2})$ $\mathcal{O}(m|v|+m^{2})$ $\mathcal{O}(m+|v|)$ $\mathcal{O}(d+\delta\cdot S\cdot m+(1-\delta)m)$ $\mathcal{O}(m\cdot|v|+m\cdot d+m\cdot S))$
e-SeaFL (Behnia et al., 2024) $\mathcal{O}(d\cdot|v|+|v|)$ $\mathcal{O}(m\cdot|v|+m)$ $\mathcal{O}(m\cdot|v|+d\cdot|v|)$ $\mathcal{O}(d\cdot|v|)$ $\mathcal{O}(m\cdot|v|)$ $\mathcal{O}(m\cdot|v|+d\cdot|v|)$
Our protocol $\mathcal{O}(d\cdot|v|+|v|)$ $\mathcal{O}(m\cdot|v|)$ $\mathcal{O}(m\cdot|v|+d\cdot|v|)$ $\mathcal{O}(d\cdot|v|)$ $\mathcal{O}(m\cdot|v|)$ $\mathcal{O}(m\cdot|v|+d\cdot|v|)$
@@
The set of decryptors in Flamingo and assistant nodes in e-SeaFL represented as the intermediate servers for comparison; $m$ represents total number of participating users; $|v|$ represents the size of the input vector; NA: Not applicable; $\sigma$ represents a security parameter and $\sigma\cdot m\leq\frac{m}{3}$; $|i|$ represents the total number of rounds in the training phase; $d$ represents the number of decryptors or intermediate servers; $\delta$ represents user dropout rate; $S$ represents the upper bound on the number of neighbors of a user; $\eta_{d}$ represents the number of corrupted decryptors; $\delta_{d}$ represents decryptor dropout rate; $\epsilon$ represents the graph generation parameter in (Ma et al., 2023).
2. Related Work
@@
Secure aggregation in FL protects individual users’ local model parameters from being disclosed to the central aggregator server. This mechanism is also referred to as privacy-preserving aggregation (Liu et al., 2022). Differential privacy (DP), homomorphic encryption (HE), secure multi-party computation (SMC), and masking methods are being used for privacy-preserving secure aggregation in FL (Yin et al., 2021). Our proposed protocol falls within the masking-based method. Therefore, we briefly introduce the other secure aggregation methods while discussing masking-based approach.
2.1. DP, HE, and SMC Based Protocols